OpenStudy (anonymous):

find the slope of the curve at y=1/x-1

OpenStudy (anonymous):

What's the curve?

OpenStudy (anonymous):

It doesn't give it, you'd have to graph it

OpenStudy (anonymous):

i'm really confused..

OpenStudy (anonymous):

If this is calculus, it is probably referring to the tangent curve...

OpenStudy (anonymous):

In which case it would be the derivative of the function

OpenStudy (anonymous):

this is most definitely calculus, but I plugged it in to the f(x+h) - f(x) all over h formula

OpenStudy (anonymous):

and ended up with 1/h^2 and I'm really confused..

OpenStudy (anonymous):

You haven't learned the derivative formulas yet??? It is much shorter

OpenStudy (anonymous):

what derivative formulas?!

OpenStudy (anonymous):

fml

OpenStudy (anonymous):

ok so ther derivative of 1/x= x^-1=x^-2 = 1/x^2. The derivative is one over x squared which is the slope

OpenStudy (anonymous):

how does x^-1 = x^-2? what?

OpenStudy (anonymous):

That is from the formula for derivatives of x to a power. you -1 from the power which in this case makes it x^-2

OpenStudy (anonymous):

I don't see how it makes it x^-2 and in the back of the book the answer says -1/(x-1)^2

OpenStudy (anonymous):

Lol. If you haven't learned the derivative formulas you have to do it the hard way and use the definition.

OpenStudy (anonymous):

could you please show me how to do it the definition way?

OpenStudy (anonymous):

or if i show you my work will you tell me what I did incorrectly?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

that is really what it says in the back of the book?!

OpenStudy (anonymous):

yes! i've tripple checked, i'm assuming it's wrong

OpenStudy (anonymous):

no wonder people get confused \[-\frac{1}{(x-1)^2}\] is not the slope of anything

OpenStudy (anonymous):

it is a FORMULA for the slope of the line tangent to the curve you need to know \(x\) to find an actual slope

OpenStudy (anonymous):

they don't ask for that though, so no idea why it's in the back of the book.. jvioejiwoeajfiewao

OpenStudy (anonymous):

it is not wrong in the sense that \(-\frac{1}{(x-1)^2}\) is not the formula for the slope, and we can find that if you like

OpenStudy (anonymous):

can you check my work for me? i'm using the derivative function because we havn't learned any formulas

OpenStudy (anonymous):

so you are working directly from the definition, right?

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\] is what i assume you are using

OpenStudy (anonymous):

yes! and it ended up like this..

OpenStudy (anonymous):

a word of advice before you start 1) is a only algebra, but you have to do it carefully 2) work step by step, and ignore the \(h\) in the denominator until all the algebra is done in the numerator

OpenStudy (anonymous):

 but then i simplified the numerator